OpenStudy (anonymous):

What are the dimensions of a rectangle that has a perimeter of 56 units and an area of 96 square units?

OpenStudy (campbell_st):

so using this given information and letting the length by l, and width w 56 = 2l + 2w making w the subject w = 1/2(56 - 2l) or w = 28 - l (equation 1) now the area 96 = lw substitute equation 1 into the area formula 96 =l(28 -l) you need to distribute and have to solve a quadratic. Hope this makes some sense.

OpenStudy (anonymous):

Thanks